# Strict Subgoal Execution: Reliable Long-Horizon Planning in Hierarchical Reinforcement Learning

### v1

@@ 5.2 Further Analysis and Ablation Studies
Fig. 8 presents a trajectory analysis of the proposed SSE framework in the AntDoubleKeyChest environment, illustrating how the agent progressively explores the map and collects both keys and the final goal. In the early stage (a) ($t\approx 300\text{K}$), the agent expands map coverage by sampling diverse subgoals, similar to simpler environments. As training progresses, the agent visits increasingly more regions, allowing it to reach the first key within a single high-level step, as shown in (b) ($t\approx 1\text{M}$). In a more advanced stage (c) ($t\approx 1.5\text{M}$), it collects both keys in just two high-level steps. Eventually, as shown in (d) ($t\approx 3\text{M}$), the agent completes the entire task, including both keys and the goal, in only three high-level steps. These results demonstrate that SSE allows the agent to reach any location in the map with a single high-level decision. As a result, it can solve complex multi-goal tasks using a minimal number of high-level steps, which highlights the effectiveness of SSE in long-horizon environments that require sequential decision-making for the high-level policy.
@@
We conduct an ablation study on AntDoubleKeyChest to evaluate the contribution of each component in SSE and the impact of key hyperparameters, as shown in Fig. 8. For the component analysis, we consider three variants: (1) SSE without path refinement (PR), which disables the failure-aware cost adjustment in graph planning; (2) SSE without $\tilde{g}_{\mathrm{novel}}$, $\tilde{g}_{\max}$, $g$, or $\pi^{\mathrm{exp}}$, where we remove each subgoal type individually or disable the exploration policy altogether; and (3) SSE with a fixed low-level step (FLS), which replaces strict subgoal enforcement with a fixed-step high-level policy as in prior HRL methods. All variants lead to degraded performance, with the fixed-step setting failing to learn entirely, emphasizing the importance of strict subgoal execution and exploratory diversity. For hyperparameter sensitivity, setting $c_{\mathrm{dist}}=5$ achieves the optimal balance between penalizing failure-prone paths and avoiding excessive detours. Smaller values have little effect, while larger ones result in inefficient routing. Grid resolution remains stable, but overly fine grids (e.g., $d_{\mathcal{G}}=1$) slow learning due to increased planning costs. For the exploration ratio $\eta$, a value of 0.2 proves optimal, as lower values limit coverage too early and higher values delay the convergence of $\pi^{h}$. Despite some sensitivity, SSE consistently outperforms all baselines, demonstrating the SSE’s superiority. Additional trajectory analyses and ablation studies in other key environments are presented in Appendix E, further demonstrating the superiority of the proposed method.
@@ 6 Limitation
While SSE demonstrates strong performance across various long-horizon goal-conditioned tasks, it is currently limited to settings where the agent’s objective is solely to reach a specified goal. This assumption restricts its applicability to more complex tasks that require sequential or composite actions, such as robotic manipulation involving grasping and relocating objects, where multiple skill primitives may be necessary. Addressing such tasks would require extending the current framework to support a broader set of low-level skills and a high-level policy capable of selecting among them, presenting a promising direction for future work. Additionally, SSE relies on discretizing the goal space into grid cells to perform failure-aware path refinement. This approach is practical for environments with low-dimensional goal spaces such as 2D or 3D, commonly found in navigation and locomotion tasks. However, in high-dimensional goal spaces, the number of partitions increases rapidly, making exploration inefficient and coverage more difficult. To address this limitation, more advanced partitioning techniques such as clustering based on learned goal representations or adaptive sampling strategies could be considered to improve scalability in complex domains.
@@ 7 Conclusion
In this paper, We proposed Strict Subgoal Execution (SSE), a graph-based HRL framework designed to improve reliability and efficiency in long-horizon, goal-conditioned tasks. By enforcing single-step subgoal reachability, SSE enables robust high-level planning and significantly reduces decision horizons. The introduction of failure-aware path refinement and a decoupled exploration policy further enhances subgoal reliability and map coverage. Extensive experiments demonstrate that SSE consistently outperforms existing HRL and graph-based methods across a range of complex tasks. These results highlight the framework’s effectiveness in enabling stable and generalizable behavior, making it a promising approach for scalable hierarchical control in various long-horizon tasks.
Appendix A Broader Impact
The proposed SSE framework enhances the practicality and robustness of HRL in sparse-reward, long-horizon environments by eliminating fixed subgoal intervals and integrating failure-aware planning. This design reduces the dependency on reward shaping and broadens the applicability of HRL to real-world domains such as robotics, autonomous navigation, and long-term task planning. Furthermore, the graph-based structure of SSE improves interpretability, making it suitable for systems demanding transparency and reliability. However, its effectiveness still depends on careful design and validation; in safety-critical applications, poorly defined goals or sudden environmental shifts could lead to suboptimal behaviors. Nevertheless, SSE does not introduce specific societal or ethical concerns beyond standard RL considerations and poses no additional risks when applied responsibly.
@@ B.1 Graph Construction of SSE
In this section, we describe how existing methods define the graph $G=(V,E)$ with the landmark node set $V$ and the edge set $E$ with edge distances $d$, and how the proposed SSE framework constructs this graph.
@@
Existing HRL methods such as DHRL [19] and NGTE [20] construct the landmark node set $V$ by selecting visited states during exploration. They employ the Farthest Point Sampling (FPS) algorithm to identify landmark nodes that are far apart from each other. This approach ensures that frequently visited regions are well-represented, as it builds $V$ based on actual agent trajectories. However, it is limited to visited states, meaning unexplored or infrequently visited areas cannot be selected as landmarks, slowing exploration in those regions. In contrast, BEAG [15] accelerates exploration by partitioning the goal space into a grid structure, creating landmark nodes at each grid intersection. This method allows for faster exploration by using virtual goal positions as landmarks, independent of visitation frequency. The structured grid layout enables more systematic and efficient exploration. To leverage this advantage, the proposed SSE framework adopts a grid-based strategy for landmark node selection inspired by BEAG. Given a 2D goal space of size $x\times y$ and a grid size of $d_{\mathcal{G}}$, the landmark set $V$ is constructed as follows:
@@
Given the landmark node set $V$, the edge set $E$ is defined as the collection of distances $d(v_{1}\rightarrow v_{2})$ between any two nodes $v_{1},v_{2}\in V$. For previous graph-based RL methods, the edge distance $d$ is computed as $d(v_{1}\rightarrow v_{2}):=\log_{\gamma^{l}}\left(1+(1-\gamma^{l})Q^{G}(v_{1},v_{2},\pi^{l})\right),$ where $\gamma^{l}$ is the low-level discount factor used for training the low-level policy $\pi^{l}$, and $Q^{G}$ is the value function estimating the traversal cost from $v_{1}$ to $v_{2}$, as described in Section 2. Existing HRL methods like DHRL [19] and NGTE [20] directly use the low-level value function $Q^{l}$, which is trained with a step-based reward of $-1$, to define the distance as the expected number of steps required for the low-level policy to navigate from $v_{1}$ to $v_{2}$. Although this method reflects actual navigation costs, it is sensitive to instability during $Q^{l}$ training, resulting in fluctuating edge distances. In contrast, BEAG [15] measures the distance using the Euclidean norm $d_{E}=||v_{1}-v_{2}||$, providing a stable but less accurate representation of traversal costs. To leverage the strengths of both approaches, SSE defines the edge distance $d(v_{1}\rightarrow v_{2})$ by combining $Q^{l}$ and $d_{E}$ as follows:
@@ B.2 Detailed Loss Functions and Implementation of the SSE Framework
As described in Section 4, the proposed SSE framework is an HRL structure that employs a high-level policy $\pi^{h}$, an exploration policy $\pi^{\textrm{exp}}$, and a low-level policy $\pi^{l}$. The exploration policy does not require separate parameterization for high-level actions, while $\pi^{h}$ and $\pi^{l}$ are parameterized by $\theta^{h}$ and $\theta^{l}$, respectively, and represented as $\pi^{h}_{\theta^{h}}$ and $\pi^{l}_{\theta^{l}}$. To evaluate these policies, SSE defines the parameterized high-level value function $Q^{h}_{\psi^{h}}$ and the low-level value function $Q^{l}_{\psi^{l}}$, where $\psi^{h}$ and $\psi^{l}$ are the respective parameters. As mentioned in Section 2, the high-level policy and value function are trained to maximize external rewards, while the low-level policy and value function are optimized to reach designated waypoints incrementally. To facilitate this, the low-level policy receives the following reward $r^{l}_{t}$ at each time step:
$r^{l}_{t}=\begin{cases}0&\text{if }\|\phi(s_{t+1})-\mathrm{wp}_{i}\|<0.5,\textrm{ (agent reaches the current waypoint)}\\ -1&\text{otherwise},\end{cases}$ (8)
where $\mathrm{wp}_{i}$ is the target waypoint at the current timestep $t$, and $(\mathrm{wp}_{1},\cdots,\mathrm{wp}_{n})$ represent the shortest path from $\phi(s_{t})$ to the subgoal $\tilde{g}_{t}$. The high-level and low-level policies, along with their value functions, are trained using the transitions stored in the high-level buffer $\mathcal{B}^{h}$ and the low-level buffer $\mathcal{B}^{l}$ through the TD3 algorithm [44], a standard off-policy RL method. The value function losses for high-level and low-level policies are defined as follows:
$\small\mathcal{L}_{Q^{h}}(\psi^{h})$ $=\mathbb{E}_{B^{h}}\left[\left(Q^{h}(\tilde{s}_{t},\tilde{g}_{t})-\left(r^{h}_{t}+\gamma^{h}\min_{i=1,2}Q^{h}_{\bar{\psi}^{h}_{i}}(\tilde{s}_{t+k_{t}},\pi^{h}_{\theta^{h}}(\tilde{s}_{t+k_{t}},g))\right)\right)^{2}\right]$
@@
where $r^{h}_{t}=\sum_{j=t}^{t+k_{t}-1}r_{j}$ for a successful trajectory and $r^{h}_{t}=0$ for a failed trajectory, $\bar{\psi}^{h}$ and $\bar{\psi}^{l}$ are target network parameters updated via exponential moving average (EMA), and $\gamma^{h}$ and $\gamma^{l}$ are the discount factors for training the high-level and low-level policies, respectively. The actor losses for optimizing the policies are defined as follows:
$\mathcal{L}_{\pi^{h}}(\theta^{h})=-\mathbb{E}_{B^{h}}\left[Q^{h}_{\psi^{h}}(\tilde{s}_{t},\pi_{\theta^{h}}^{h}(\tilde{s}_{t},g))\right],\quad\mathcal{L}_{\pi^{l}}(\theta^{l})=-\mathbb{E}_{B^{l}}\left[Q^{l}_{\psi^{l}}(s_{t},\pi_{\theta^{l}}^{l}(s_{t},\mathrm{wp}_{i}))\right].$ (10)
@@
The parameters are optimized using the Adam optimizer [45] to minimize the respective loss functions. SSE distinguishes between the high-level discount factor $\gamma^{h}$ and the low-level discount factor $\gamma^{l}$, setting $\gamma^{l}=0.99$ as typical in RL, and $\gamma^{h}=0.4$ to limit return propagation across high-level steps, encouraging shorter path optimization. In the initial stages of training, rewards are sparse. To address this, the high-level buffer $\mathcal{B}^{h}$ is divided evenly, with half storing successful trajectories and the other half storing trajectories with zero reward. This design improves learning signals from successful experiences. The exploration ratio, which controls the balance between the exploration policy and the high-level policy, starts at $1{:}0$ and decays by 0.05 per iteration until it reaches $\eta{:}(1{-}\eta)$, as outlined in Section 4. This scheduling promotes exploration initially and shifts the focus to high-level learning as training progresses. If the low-level agent fails to reach the high-level subgoal $\tilde{g}_{t}$, the trajectory is marked as failed. In cases where the agent becomes stuck, such as flipping over or hitting obstacles, its position may remain unchanged for long periods. To improve sample efficiency, if no movement is detected for 500 steps, the trajectory is classified as failed and the episode is terminated. The complete SSE algorithm is provided in Algorithm 2.
@@
Initialize: Policies $\pi^{h}$, $\pi^{\mathrm{exp}}$, $\pi^{l}$, buffers $B^{h}$, $B^{l}$, grid cells $C_{\mathcal{G}}$
@@
                   Sample a subgoal $\tilde{g}_{t}\sim\pi(\tilde{s}_{t},g)$
@@
                         Roll out the low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ to reach the waypoint $\mathrm{wp}_{i}$ within $k_{t}$ steps
                        Store the $k_{t}$ transitions $(s_{t},\mathrm{wp}_{i},a_{t},r_{t},s_{t+1})$ into the low-level buffer $\mathcal{B}^{l}$
                        Compute the reward sum: $r_{\mathrm{sum}}=\sum_{j=t}^{t+k_{t}-1}r_{j}$
@@
                               Success: Store the high-level transition $(\tilde{s},g,\tilde{g}_{t},r_{\mathrm{sum}},\tilde{s}_{t+k_{t}})$ into $B^{h}$
@@
                               Failure: Store $(\tilde{s},g,\tilde{g}_{t},0,\tilde{s}_{T})$ into $B^{h}$
                              $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})+=1$ for $m$ s.t. $\phi(s_{t+k_{t}})\in C_{\mathcal{G}}^{m}$
@@
      Update $\psi^{h}$, $\psi^{l}$, $\theta^{h}$, $\theta^{l}$ to minimize RL losses $\mathcal{L}_{Q^{h}}$, $\mathcal{L}_{Q^{l}}$, $\mathcal{L}_{\pi^{h}}$, $\mathcal{L}_{\pi^{l}}$, respectively

### v2

@@ 5.2 Further Analysis and Ablation Studies
In the following analysis and ablation studies, we focus on the grid-based approach and refer to SSE (Grid) simply as SSE, as it consistently achieves better performance and offers a richer set of hyperparameters for analysis and implementation. Fig. 7 presents a trajectory analysis of the proposed SSE framework in the AntDoubleKeyChest environment, illustrating how the agent progressively explores the map and collects both keys and the final goal. In the early stage (a) ($t\approx 300\text{K}$), the agent expands map coverage by sampling diverse subgoals, similar to simpler environments. As training progresses, the agent visits increasingly more regions, allowing it to reach the first key within a single high-level step, as shown in (b) ($t\approx 1\text{M}$). In a more advanced stage (c) ($t\approx 1.5\text{M}$), it collects both keys in just two high-level steps. Eventually, as shown in (d) ($t\approx 3\text{M}$), the agent completes the entire task, including both keys and the goal, in only three high-level steps. These results demonstrate that SSE allows the agent to reach any location in the map with a single high-level decision. As a result, it can solve complex multi-goal tasks using a minimal number of high-level steps, which highlights the effectiveness of SSE in long-horizon environments that require sequential decision-making for the high-level policy. Notably, SSE solves this sequential task without an explicit curriculum. The reliability of its high-level policy, learned via FER, combined with an augmented state including key-possession flags, enables the agent to autonomously discover the required sequence of sub-objectives.
@@
We conduct an ablation study on AntDoubleKeyChest to evaluate the contribution of each component in SSE and to analyze the impact of key hyperparameters, as shown in Fig. 8. For the component analysis, we consider five variants: (1) SSE with FPS, which replaces the grid-based landmarks with Farthest Point Sampling Kim et al. (2021); (2) SSE with HER, which replaces strict subgoal execution with a conventional high-level policy that uses HER; (3) SSE without $\mathcal{B}_{F}^{h}$, which disables FER by using a standard replay buffer; (4) SSE without $\pi^{\mathrm{exp}}$, which disables the exploration policy; and (5) SSE without path refinement (PR), which disables the failure-aware adjustment. All variants exhibit degraded performance. Notably, the settings using HER or disabling FER fail to learn entirely, emphasizing the importance of strict subgoal execution and the consistent high-level signals provided by FER. SSE with FPS eventually succeeds, though less efficiently than the grid version, while still outperforming FPS-based baselines. This confirms that our framework remains effective regardless of the specific landmark selection strategy. Our hyperparameter analysis shows that setting $c_{\mathrm{dist}}=5$ achieves an optimal balance, though performance holds steady across a range of values. Similarly, performance also holds steady for grid resolution, though overly fine grids (e.g., $d_{\mathcal{G}}=1$) can slow learning. For the exploration ratio $\eta$, a value of 0.2 proves optimal. While hyperparameter tuning optimizes performance, SSE consistently outperforms all baselines. Additional analyses in other environments and sensitivity studies on the reachability threshold $\lambda$ and the high-level discount factor $\gamma^{h}$ are provided in Appendix E, further characterizing SSE.
@@ 6 Limitation
Our framework introduces new hyperparameters, such as the exploration ratio $\eta$ and path refinement factor $c_{\mathrm{dist}}$, which require tuning. However, we find their effective ranges to be stable across diverse environments, and our ablation studies confirm that the framework maintains strong performance across a wide range of values, minimizing the overall tuning cost. While SSE also adds computational steps, its principle of early termination on subgoal failure yields significant efficiency gains. This prevents long, unproductive trajectories and results in a faster computation time compared to other recent methods, as quantitatively analyzed in Appendix D.2.
@@ 7 Conclusion
In this paper, we proposed SSE, a graph-based HRL framework designed to improve reliability and efficiency in long-horizon, goal-conditioned tasks. By enforcing single-step subgoal reachability, SSE enables more direct and reliable high-level planning and significantly reduces decision horizons. The introduction of failure-aware path refinement and a decoupled exploration policy further enhances subgoal reliability and map coverage. Extensive experiments demonstrate that SSE consistently outperforms existing HRL and graph-based methods across a range of complex tasks. These results highlight the framework’s effectiveness in enabling stable and generalizable behavior, making it a promising approach for scalable hierarchical control in various long-horizon tasks.
Acknowledgment
This work was supported partly by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-II220469, Development of Core Technologies for Task-oriented Reinforcement Learning for Commercialization of Autonomous Drones), (No. RS-2025-25442824, AI Star-Fellowship Program (UNIST)), and (No. RS-2020-II201336, Artificial Intelligence Graduate School Support (UNIST)), and partly by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2025-23523191, LLM-Based Multi-Agent Reinforcement Learning for End-to-End Large Autonomous Swarm Control).
Ethics statement
This paper introduces a foundational algorithm, SSE, designed to improve the long-horizon planning capabilities of reinforcement learning agents. All experiments were conducted in standard, simulated robotics environments. As such, our research does not involve human subjects, sensitive or personally identifiable data, nor does it directly address systems that interact with people. Therefore, issues of data privacy, dataset bias, and fairness are not directly applicable to this work.
Reproducibility statement
We are committed to the reproducibility of our research. The complete source code for our proposed framework, SSE, and all experiments is included as an anonymized zip file in the supplementary materials. A detailed breakdown of the implementation, including network architectures and training procedures, can be found in Appendix B. All hyperparameters required to reproduce our results are provided in Appendix C.3, with common settings listed in Table 2 and environment-specific configurations in Table 3. Furthermore, Appendix C details the full experimental setup, including descriptions of the baseline algorithms, specifications for all environments (Table 1), and the hardware and software configurations on which the experiments were conducted. We believe these resources will enable the reproduction of our findings.
Appendix A The Use of Large Language Models
We utilized a large language model (LLM) as an assistive tool during the preparation of this manuscript. The LLM’s role was strictly limited to polishing the text, which includes improving clarity, conciseness, and correcting grammatical errors. The LLM was not used for research ideation. The authors have carefully reviewed and edited all content and take full responsibility for the scientific accuracy and integrity of this work. The LLM is not credited as an author.
@@ B.1 Graph Construction of SSE
In this section, we describe how existing methods define the graph $G=(V,E)$,which consists of a landmark node set $V$ and the edge set $E$ with edge distances $d$. We then explain how the proposed SSE framework constructs this graph.
@@
Existing HRL methods such as DHRL (Lee et al., 2022) and NGTE (Park et al., 2024) construct the landmark node set $V$ by selecting visited states during exploration. They employ the Farthest Point Sampling (FPS) algorithm to identify landmark nodes that are far apart from each other. This approach ensures that frequently visited regions are well-represented, as it builds $V$ based on actual agent trajectories. However, it is limited to visited states, meaning unexplored or infrequently visited areas cannot be selected as landmarks, slowing exploration in those regions. In contrast, BEAG (Yoon et al., 2024) accelerates exploration by partitioning the goal space into a grid structure, creating landmark nodes at each grid intersection. This method allows for faster exploration by using virtual goal positions as landmarks, independent of visitation frequency. The structured grid layout enables more systematic and efficient exploration. To leverage this advantage and ensure a structured, reproducible setup, SSE adopts a grid-based landmark selection strategy inspired by BEAG. Given a 2D goal space of size $x\times y$ and a grid size of $d_{\mathcal{G}}$, the landmark set $V$ is constructed as follows:
@@
[…]
@@ B.2 Detailed Loss Functions and Implementation of the SSE Framework
RL training losses
As described in Section 4, the proposed SSE framework is an HRL structure that employs a high-level policy $\pi^{h}$, an exploration policy $\pi^{\textrm{exp}}$, and a low-level policy $\pi^{l}$. The exploration policy does not require separate parameterization for high-level actions, whereas $\pi^{h}$ and $\pi^{l}$ are parameterized by $\theta^{h}$ and $\theta^{l}$, respectively, and are represented as $\pi^{h}_{\theta^{h}}$ and $\pi^{l}_{\theta^{l}}$. To evaluate these policies, SSE defines the parameterized high-level value function $Q^{h}_{\psi^{h}}$ and the low-level value function $Q^{l}_{\psi^{l}}$, where $\psi^{h}$ and $\psi^{l}$ are the respective parameters. As mentioned in Section 2, the high-level policy and value function are trained to maximize external rewards, while the low-level policy and value function are optimized to reach designated waypoints incrementally. To facilitate this, the low-level policy receives the reward $r^{l}_{t}$ at each time step, defined as follows:
$r^{l}_{t}=\begin{cases}0&\text{if }\|\phi(s_{t+1})-\mathrm{wp}_{i}\|<\lambda,\textrm{ (agent reaches the current waypoint)}\\ -1&\text{otherwise},\end{cases}$ (8)
where $\lambda$ is the reachability check threshold, $\mathrm{wp}_{i}$ is the target waypoint at the current timestep $t$, and $(\mathrm{wp}_{1},\cdots,\mathrm{wp}_{n})$ represent the shortest path from $\phi(s_{t})$ to the subgoal $\tilde{g}_{t}$. The high-level and low-level policies, along with their value functions, are trained using the transitions stored in the FER $\mathcal{B}_{F}^{h}$ and the low-level buffer $\mathcal{B}^{l}$ through the TD3 algorithm (Fujimoto et al., 2018), a standard off-policy RL method. The value function losses for high-level and low-level policies are defined as follows:
$\mathcal{L}_{Q^{h}}(\psi^{h})$ $=\mathbb{E}_{B_{F}^{h}}\left[\left(Q^{h}(s_{t},\tilde{g}_{t})-\left(r^{h}_{t}+\gamma^{h}\min_{i=1,2}Q^{h}_{\bar{\psi}^{h}_{i}}(s_{t^{\prime}},\pi^{h}_{\theta^{h}}(s_{t^{\prime}},g))\right)\right)^{2}\right]$
@@
where $t^{\prime}$ denotes the termination time of the low-level path execution, which is variable as the step concludes only upon success or failure. In the case of a partial success, $t_{\mathrm{wp}}$ (used to store transitions in FER) indicates the time step when the agent reached the last successful waypoint $wp_{\mathrm{final}}$. The high-level reward is then defined as $r^{h}_{t}=\sum_{j=t}^{t^{\prime}-1}r_{j}$ for a successful trajectory and $r^{h}_{t}=0$ for a failed one. The terms $\bar{\psi}^{h}$ and $\bar{\psi}^{l}$ are target network parameters updated via exponential moving average (EMA), and $\gamma^{h}$ and $\gamma^{l}$ are the discount factors for training the high-level and low-level policies, respectively. The actor losses for optimizing the policies are defined as follows:
$\mathcal{L}_{\pi^{h}}(\theta^{h})=-\mathbb{E}_{B_{F}^{h}}\left[Q^{h}_{\psi^{h}}(s_{t},\pi_{\theta^{h}}^{h}(s_{t},g))\right],\quad\mathcal{L}_{\pi^{l}}(\theta^{l})=-\mathbb{E}_{B^{l}}\left[Q^{l}_{\psi^{l}}(s_{t},\pi_{\theta^{l}}^{l}(s_{t},\mathrm{wp}_{i}))\right].$ (10)
@@
The parameters are optimized using the Adam optimizer (Kingma, 2014) to minimize the respective loss functions. SSE distinguishes between the high-level discount factor $\gamma^{h}$ and the low-level discount factor $\gamma^{l}$, setting $\gamma^{l}=0.99$ as typical in RL, and $\gamma^{h}=0.4$ to limit return propagation across high-level steps, encouraging shorter path optimization. In the initial stages of training, rewards are sparse. To address this, the FER $\mathcal{B}_{F}^{h}$ is divided evenly, with half storing successful trajectories and the other half storing trajectories with zero reward. This design improves learning signals from successful experiences. The exploration ratio, which controls the balance between the exploration policy and the high-level policy, starts at 1:01{:}0 and decays by 0.05 per iteration until it reaches $\eta{:}(1{-}\eta)$, as outlined in Section 4. This scheduling promotes exploration initially and shifts the focus to high-level learning as training progresses. If the low-level agent fails to reach the high-level subgoal $\tilde{g}_{t}$, the trajectory is marked as failed. In cases where the agent becomes stuck, such as flipping over or hitting obstacles, its position may remain unchanged for long periods. To improve sample efficiency, if no movement is detected for 500 steps, the trajectory is classified as failed and the episode is terminated. The complete SSE algorithm is provided in Algorithm 2.
Model training losses for SSE (Model)
We now describe the implementation details of the prediction network $f_{\zeta}$, which computes the prediction error used to determine the novel region in the model-based approach, and the failure predictor network $F_{\xi}$, which estimates the failure ratio of target nodes.
In the model-based variant of the exploration policy, we estimate the novel region $V_{\mathrm{novel}}^{\mathrm{model}}$ using Random Network Distillation (RND) (Burda et al., 2018), which was originally proposed in standard RL to identify novel states based on prediction error. Following the RND formulation, we use a fixed target network $f_{\zeta_{\mathrm{targ}}}:V\rightarrow\mathbb{R}^{k}$ and a predictor network $f_{\zeta}:V\rightarrow\mathbb{R}^{k}$ with parameters $\zeta$. The predictor $f_{\zeta}$ is trained on visited sample nodes stored in the buffer $\mathcal{D}$ to regress the target output via the mean squared error objective
$L_{\mathrm{RND}}(\zeta)=\mathbb{E}_{v\sim\mathcal{D}_{\mathrm{visited}}}\left[\left\|f_{\zeta}(v)-f_{\zeta_{\mathrm{targ}}}(v)\right\|_{2}^{2}\right].$ (11)
As the predictor network is trained to minimize the prediction error on visited nodes, the prediction error becomes small in frequently explored regions and remains large for underexplored nodes. We therefore use the squared prediction error $\left\|f_{\zeta}(v)-f_{\zeta_{\mathrm{targ}}}(v)\right\|_{2}^{2}$ as a novelty score, and nodes with larger error are treated as more novel. By selecting nodes with high prediction error as exploration targets, the model-based variant is encouraged to visit underexplored regions of the goal space.
For failure-aware refinement, we train a failure predictor network $F_{\xi}:V\rightarrow[0,1]$ parameterized by $\xi$, which outputs the estimated failure probability of each node. The network is trained using node samples stored in buffer $\mathcal{D}_{v}$, where each node $v$ is paired with a binary label $y\in\{0,1\}$ indicating whether the trajectory containing $v$ failed $(y=1)$ or succeeded $(y=0)$. We optimize $F_{\xi}$ with the binary cross-entropy loss
$L_{F}(\xi)=-\mathbb{E}_{(v,y)\sim\mathcal{D}_{v}}\left[y\log(F_{\xi}(v))+(1-y)\log(1-F_{\xi}(v))\right].$ (12)
Through this training, $F_{\xi}(v)$ approximates the empirical failure ratio at node $v$, which we use as the model-based failure estimate in our refinement scheme.
@@
Initialize: Policies $\pi^{h}$, $\pi^{\mathrm{exp}}$, $\pi^{l}$, buffers $B_{F}^{h}$, $B^{l}$, grid cells $C_{\mathcal{G}}$ (Grid) or Models $f_{\zeta},F_{\xi}$ (Model)
@@
         Sample a subgoal $\tilde{g}_{t}\sim\pi(s_{t},g)$
@@
            Roll out the low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ to reach the waypoint $\mathrm{wp}_{i}$
           Store the $t^{\prime}-t$ transitions $(s_{t},\mathrm{wp}_{i},a_{t},r_{t},s_{t+1})$ into the low-level buffer $\mathcal{B}^{l}$
           Compute the reward sum: $r_{t}^{h}=\sum_{j=t}^{t^{\prime}-1}r_{j}$
          Construct the FER:
@@
             Success: Store the success transition $(s,g,\tilde{g}_{t},r_{\mathrm{sum}},s_{t^{\prime}})$ into FER $B_{F}^{h}$
@@
             Failure: Store the stop-on-failure transition $(s,g,\tilde{g}_{t},0,s_{T})$ into FER $B_{F}^{h}$
              Store the partial success transition $(s,g,\mathrm{wp}_{\mathrm{final}},\sum_{j=t}^{t_{\mathrm{wp}}-1}r_{j},s_{t_{\mathrm{wp}}})$ into FER $B_{F}^{h}$
@@
  Update total visitation count $N(C_{\mathcal{G}}^{m})$ and failure count $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})$ for all $C_{\mathcal{G}}^{m}$ (Grid-based)
  Update model parameters $\zeta$, $\xi$ to minimize $L_{\mathrm{RND}}(\zeta)$ and $L_{\mathrm{F}}(\xi)$ (Model-based)
  Update $\psi^{h},\theta^{h}$ using samples from FER $B_{F}^{h}$ to minimize $\mathcal{L}_{Q^{h}}(\psi^{h}),\mathcal{L}_{\pi^{h}}(\theta^{h})$
  Update $\psi^{l},\theta^{l}$ using samples from $\mathcal{B}^{l}$ to minimize $\mathcal{L}_{Q^{l}}(\psi^{l}),\mathcal{L}_{\pi^{l}}(\theta^{l})$